# 5- and 6-membered rings: A natural orbital functional study

### v1

@@ V Closing remarks
This study assesses the performance of the most recent electron-pairing-based natural orbital functionals, GNOF and its modified variant GNOFm, on absolute correlation energies for five- and six-membered rings. This benchmark set, composed of simple aromatic rings of broad relevance, has previously been used to examine the performance and convergence properties of the Møller–Plesset series and coupled-cluster methods (including iterative approximations). Our results show that GNOFm attains quantitative agreement with the ground-state reference CCSD(T) across multiple sizes of the correlation-consistent Dunning basis sets. We also report complete-basis-set (CBS) extrapolated correlation energies for GNOF, GNOFm, and CCSD(T). A direct comparison between GNOFm and CCSD(T) indicates agreement to approximately 50 mEh, suggesting that the present data can serve as a useful benchmark for other quantum-chemistry approaches.
@@
The benchmarking dataset used here will be extended in a forthcoming work to guide refinements of the functional form, with particular emphasis on improving the treatment of electron correlation within the NOFT framework. Overall, the findings support the continued development of NOFs as a viable alternative to traditional density functionals and multireference wave-function methods. Their ability to capture both static and dynamic correlation without active-space selection makes them especially attractive for complex chemical systems, particularly when large molecules are involved. Future efforts should focus on further improving the balanced description of dynamic and non-dynamic correlations within NOFT to enhance accuracy across a broader range of chemical environments.

### v2

@@ V Closing remarks
This study assesses the performance of the most recent electron-pairing-based natural orbital functionals, GNOF and its modified variant GNOFm, on absolute correlation energies for five- and six-membered rings. This benchmark set, composed of simple aromatic rings of broad relevance, has previously been used to examine the performance and convergence properties of the Møller–Plesset series and coupled-cluster methods (including iterative approximations). Our results show that GNOFm attains qualitative agreement with the ground-state reference CCSD(T) across multiple sizes of Dunning’s weighted core-valence cc-pwCVNZ basis sets. We also report complete-basis-set (CBS) extrapolated correlation energies for GNOF, GNOFm, and CCSD(T). A direct comparison between GNOFm and CCSD(T) indicates agreement to approximately 50 mEh, suggesting that GNOFm can be employed to describe systems with dynamic electron correlation effects, in contrast to previous electron-pairing based NOFs.
@@
The benchmarking dataset used here will be extended in a forthcoming work to guide refinements of the functional form, with particular emphasis on improving the treatment of electron correlation within the NOFT framework. Overall, the findings support the continued development of NOFs as a viable alternative to traditional density functionals and multireference wavefunction methods. Their ability to capture both static and dynamic correlation without active-space selection makes them especially attractive for complex chemical systems, particularly when large molecules are involved. Future efforts should focus on further improving the balanced description of dynamic and non-dynamic correlations within NOFT to enhance accuracy across a broader range of chemical environments. In addition, and following the dipole moments reported in Table 4, our laboratory is investigating the impact of incorporating dynamic correlation effects directly into the functional on the charge distribution. The latter implies not only electric multipole moments such as dipoles or quadruples, but also equilibrium geometries.
Data availability
The materials that support the findings of this study are openly available at https://osf.io/64pcn/overview?view_only=72e6a504f07042bbb6d7ccf55aaf29c6.
Supplementary Information
Additional calculations carried out with the Dunning’s cc-pVXZ (X=2-5) basis sets, as well as a detailed study of dipole moments.
@@ VI Acknowledgments
Supplementary Material
VI.1 Calculations with the Dunning’s cc-pVXZ basis set
In the present section, we show correlation energies obtained for the correlation-consistent Dunning’s basis sets (cc-pVXZ, X=2-5). The latter are preferable to be used in calculations with frozen core electrons, however, due to their relevance to the historical development of NOF approximations and their widespread use, we include the corresponding correlation energy analysis below.
Figure S1: Correlation energies ($E-E_{\mathrm{HF}}$) in mEh for the selected set of molecules, obtained by using GNOF (dashed lines), GNOFm (dotted lines), and CCSD(T) (solid lines) using the cc-pVXZ basis sets. The cardinal numbers X = 2, 3, 4, and 5 correspond to blue, red, green and orange, respectively. Molecules ordered according to the numbering given in Table 3.
Correlation energies for GNOF, GNOFm, and CCSD(T) are shown in Fig. S1 for increasing size cc-pVXZ basis sets, being X=2-5 the cardinal number of the basis. Here, molecules are ordered according to the numbering presented in Table 3. NOF and CCSD(T) curves are roughly parallel, so the molecular description agrees for both methods independently of the size of the basis set, as well as of the different studied molecular rings.
No. Systems GNOF GNOFm CCSD(T)
1 Cyclopentadiene -1051.3 -1084.4 -1110.6
2 Pyrrole -1081.7 -1118.9 -1149.0
3 Furan -1114.7 -1148.2 -1173.7
4 Imidazole -1118.3 -1153.1 -1183.4
5 Benzene -1245.5 -1265.8 -1304.0
6 Pyridine -1262.8 -1301.7 -1338.9
7 Pyrazine -1302.9 -1343.0 -1375.2
8 Pyrimidine -1298.5 -1338.3 -1371.6
9 Pyridazine -1299.9 -1346.6 -1379.0
10 Triazine -1320.3 -1374.8 -1402.0
11 Thiophene -1179.6 -1214.3 -1204.5
12 Tetrazine -1372.4 -1419.3 -1453.8
Table 3: Complete basis set (CBS) extrapolated correlation energies ($E-E_{\mathrm{HF}}$) in mEh for the 12 molecular systems, computed using GNOF, GNOFm, and CCSD(T). Helgaker’s extrapolation scheme, $E_{\infty}+bX^{-3}$, was employed with $X=2,3,4,5$ as the cardinal number of the basis set.
GNOF and GNOFm complete basis set limit (CBS) correlation energy estimates are given in Table 3, which also presents CCSD(T) values, performed using the same procedure. Helgaker’s extrapolation scheme, $E_{\infty}+bX^{-3}$ with $X=2,3,4,5$, was employed to carry out the CBS extrapolation. An inspection of CBS limit molecular correlation energies reveals an agreement within 100 mEh between GNOF and CCSD(T), values that are even improved to around 50 mEh when GNOFm is utilized. The results shown in Table 3 are summarized in Fig. S2. GNOFm energies systematically get closer to CCSD(T) results when the static term between electron pairs is modified as described in the section II from the manuscript. In other words, Fig. S2 reveals that GNOFm CBS correlation energies reduce differences between GNOF and CCSD(T) to the half. Additionally, as it is shown in Fig. S1, the improvement is obtained for all basis sets studied.
Figure S2: Complete basis set (CBS) extrapolated correlation energies ($E-E_{\mathrm{HF}}$) in mEh for the 12 molecular systems, computed using GNOF, GNOFm, and CCSD(T). Helgaker’s extrapolation scheme, $E_{\infty}+bX^{-3}$, was employed with $X=2,3,4,5$ as the cardinal number of the basis set.
VI.2 Molecular dipole moments
In this section, we show the dipole moments obtained by using GNOF and GNOFm approximations, together with reference Hartree-Fock and experimental values. Literature sources corresponding to the latter, as well as numerical values (in a.u.), are given in Table 4. Both GNOF and GNOFm improves Hartree-Fock values and corresponding dipole moments get closer to experimental data when the basis set is increased beyond cc-pwCVDZ, as it is shown for Pyrrole, Imidazole, or Pyridazine. GNOFm provides the most accurate values for the largest cc-pwCVQZ basis set, whereas GNOF dipole moments are similar, in average, for the cc-pwCVTZ and cc-pwCVQZ basis sets. Thiophene molecule corresponds to largest errors, and the latter do not improve if $d$ functions are removed from the basis set. In fact, corresponding dipole moments for the cc-pwCVQZ basis set read as 0.2703 a.u. and 0.2833 a.u., respectively for GNOF and GNOFm, which do not change significantly the errors obtained with the aug-cc-pwCVQZ basis set as it is shown in Table 4. HF gives same values regardless of the cardinal number of the basis set, as discussed in the manuscript.
Systems 1 2 3 4 6 8 9 11
HF/$c$-pwCVXZ 1.222 0.032 1.914 1.965 2.213 2.242 3.823 3.662
GNOF/$c$-pwCVDZ 0.160 0.776 0.270 1.480 0.851 0.904 1.639 0.258
GNOF/$c$-pwCVTZ 0.159 0.737 0.283 1.475 0.870 0.910 1.655 0.266
GNOF/$c$-pwCVQZ 0.165 0.727 0.291 1.476 0.885 0.913 1.663 0.271
GNOFm/$c$-pwCVDZ 0.163 0.762 0.271 1.470 0.847 0.893 1.627 0.283
GNOFm/$c$-pwCVTZ 0.159 0.731 0.282 1.463 0.856 0.895 1.646 0.270
GNOFm/$c$-pwCVQZ 0.167 0.724 0.286 1.462 0.886 0.904 1.656 0.271
Exp. 0.165 0.695 0.268 1.444 0.871 0.918 1.660 0.216
Ref. (1) (2) (3) (4) (5) (6) (7) (8)
Table 4: Non-zero dipole moments (in a.u.) computed using HF, GNOF and GNOFm with Dunning’s $c$-pwCVXZ basis sets ($X=2,3,4$), together with experimental values and their corresponding reference from the literature. Results corresponding to thiophene (no. 11) are obtained with the aug-$c$-pwCVXZ basis sets. Molecular numbering reads as follows: 1. Cyclopentadiene, 2. Pyrrole, 3. Furan, 4. Imidazole, 6. Pyridine, 8. Pyrimidine, 9. Pyridazine, 11. Thiophene.